Check the divisibility of 21436587 by 9.

djcbvweyhvb wdhnklqecnfuuiqnmckne

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The sum of the digits of 21436587 is 2 + 1 + 4 + 3 + 6 + 5 + 8 + 7 = 36.
This number is divisible by 9 (for 36 ÷ 9 = 4).

Hence, 21436587 is divisible by 9.

Check method :21436587/9 =2381843

Thanks